How many strings of 4 lower case English letters are there that have the letter x in them somewhere? Here strings may use the same letter more than once. (Hint: It might be easier to first count the strings that don't have an x in them.)
I already know that there is overall a total of 456976 total possabilities if all letters can be used and repeated but i dont know where to go from there

Answer by edjones(8007)

25^4=390625 four letter words without an x.
26^4=456976 possible 4 letter words.
456976-390625=66,351 four letter words with an x.
.
Ed